OpenStudy (brightemmerald):

The function A(b) relates the area of a trapezoid with a given height of 10 and one base length of 7 with the length of its other base. It takes as input the other base value, and returns as output the area of the trapezoid. A(b)=10*(B+7)/2 Which equation below represents the inverse function B(a), which takes the trapezoid's area as input and returns as output the length of the other base?

OpenStudy (caozeyuan):

well, how do you change the equation so that the right side is noly b

OpenStudy (caozeyuan):

A=5(B+7), B+7=A/5,

OpenStudy (brightemmerald):

Help please, at least a little bit more.

OpenStudy (3mar):

If I may help?

OpenStudy (brightemmerald):

Of course!

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

You have the equation A(b)=10*(B+7)/2, and you want to get the inverse of this?

OpenStudy (brightemmerald):

Yes

OpenStudy (3mar):

Ok. Here are the steps 1. solve for b, which means separate all b terms in one side and all other terms at the other side. 2. swap b and A(b) you get the inverse function!

OpenStudy (3mar):

it would better to show me how you did the math!

OpenStudy (3mar):

@BrightEmmerald Are you there? Can we proceed?

OpenStudy (3mar):

\[A(b)=10*\frac{ B+7 }{ 2 }=5(B+7)\] \[A(b)=5(B+7)\] \[\frac{ A(b) }{ 5 }=B+7\] \[\frac{ A(b) }{ 5 }-7=B\] now you should to swap the variables to find the inverse function: \[\frac{ B }{ 5 }-7=A(b)\] \[A(b)=\frac{ B }{ 5 }-7\] This is what you are looking for!